The supplement of an angle is 4 more than three times the angle. Find the angle. (Two angles are supplemental if their sum is 180).
Answer by cleomenius(959)

180 -x -4 = 3x
176 = 4x
44 degrees = x, the angle.
180 - 44 = 136, the supplement.
44* 3 = 132, + 4 = 136 degrees.
This does check.
Cleomenius.